### v2

@@ 3.1.1 Light curves
The ZTF light curve for NGC 6833 is shown in Figure 1 (top panel). Fortunately, this target has good -band coverage. The most prominent feature is the regular brightening events in the ZTF and the redder bands, with symmetric ingress and egress. The amplitude is in the range of , with the largest amplitude in -band. We note here though that the nebular flux can reduce the perceived variability, so the intrinsic amplitude of the source (likely the central star) can be much higher (see a brief discussion in Section 3.1.3). The less prominent but more intriguing feature appears in the ZTF band, which exhibits shallow (amplitude of ) dips nearly coincident with the peaks in the redder bands. The behavior is periodic and significant peaks appear in the Lomb-Scargle periodogram in all the three ZTF bands. The result is shown in the top panel of Figure 2. The detected periods in the three bands are marginally different (which is not surprising with long periods), with a mean value of days. The phase-folded ZTF light curve at this mean period is shown in the middle panel of the same figure. Here, the activity appears to be slightly phase-delayed from the redder to the bluer bands. However, the current data is not enough to establish this property with high confidence.
Before any further discussion, we caution that the brightness of this object ( mag in and mag in ) is very close to the saturation limit of ZTF ( in all bands). In fact, in one of the ZTF fields, the reference image is indeed saturated in the -band (where the object is brightest, mainly due to the very strong [O III] emission lines). Though such data points get rejected through the imposed quality cuts, the chance remains that the other flux readings are unreliable. To ascertain that this has not significantly affected the observed variability in the -band, especially the dip, we checked the -band light curve from the All Sky Automated Survey for SuperNovae (ASAS-SN) light curve provided at Sky Patrol V2.0333http://asas-sn.ifa.hawaii.edu/skypatrol/ (Shappee et al. 2014; Hart et al. 2023), which has a larger saturation brightness ( magnitude). The ASAS-SN -band light curve displays similar behavior as ZTF and the dips are visible. The data quality, however, is worse than ZTF thus not shown. For ZTF and bands, manual inspection shows that neither the reference nor the science images are saturated. We thus proceed with the assumption that the light curves are trustworthy.
3.1.2 Follow-up Optical Spectroscopy
NGC 6833 was observed with the DouBle SPectrograph (DBSP; Oke & Gunn 1982) attached to the Cassegrain focus of the Palomar 200-in Hale Telescope. We used the D55 dichroic, the 600 line mm-1 grating for the blue arm blazed at 3780 Å, and the 316 line mm-1 grating for the red arm blazed at 7150 Å. We used grating angles of 27∘17′ and 24∘38′ for the blue and red sides, respectively. With these setups and a slit width of 1.5′′ long-slit, we achieve resolving powers of in the blue arm and in the red arm. These setups provide continuous spectral coverage across 2900–10800 Å, with the division between blue and red arms typically occurring around 5650 Å. We followed the same reduction procedure as described in Section 4.2.1 in Paper I. Briefly, we used PypeIt (Prochaska et al., 2020a, b) with the options use_2dmodel_mask = False and no_local_sky = True to avoid masking and local subtraction of the strong emission lines. Following this, we perform manual trace identification on the two-dimensional spectrum. Figure 3 shows the optical spectrum for NGC 6833.
As mentioned previously, the object is spectroscopically well-studied (for example, Wright et al. 2005; Wesson et al. 2005; Hyung et al. 2010). However, given the knowledge of the photometric variability, the epoch of this spectrum is “special” as the object is nearly at its photometric extremum. We attempted to estimate the phase at which the previous spectroscopic observations were made, using the observation dates mentioned in the corresponding papers. Assuming that the ZTF inferred period is stable, the best estimate places the optical observations of Wright et al. (2005) and Hyung et al. (2010) at the quiescent phase. However, no significant difference in the spectral features was observed compared to the earlier studies. The calculated emission line fluxes of the major emission lines are also in very good agreement with those previously reported (in Wright et al. 2005, for example). However, a more thorough comparison is needed to identify differences at the level that can induce the observed variability (which includes ensuring consistency of trace selection, seeing, instrumental effects, etc.). This is beyond the scope of this work.
@@ 3.1.3 Discussion
The stable variability and the associated periodicity strongly suggest binarity, with the period of variability as the binary orbital period. Long-period binaries inside PNe are only rarely observed, making NGC 6833 an important addition. Only three CSPNe are known to have an orbital period longer than the inferred period of NGC 6833: BD+332642 (Van Winckel et al., 2014; Jones et al., 2017), LoTr 5, and NGC 1514 (Jones et al., 2017). The shape of the light curve and the larger amplitudes in the redder bands suggest an irradiation effect. In that case, the relatively high variability amplitude (at this period) requires the orbit to be eccentric, which is not unexpected in long-period binaries. All the other known long-period binaries in CSPNe are significantly eccentric. Eccentricity pumping mechanisms like orbital phase-dependent mass transfer can indeed maintain high eccentricities up to periods of days (see for example Bonačić Marinović et al. 2008). At this period, it is also at a margin of being a post-CE system.
Even within the binary scenario, mechanisms other than irradiation effect can possibly contribute to the observed variability. For example, the light curves in ZTF- and redder bands bear a strong resemblance to several RV Tauri (evolved pulsating giants) systems showing long-term photometric modulation (known as the RVb phenomenon, see for example the light curves in Kiss & Bódi 2017 which shows very similar ‘triangle-shaped’ brightenings with similar periods in the range of a few years). The widely accepted explanation is that the variability results from the phase-dependent dust-obscuration of the flux-dominant star in the presence of a thick circumbinary disc (Van Winckel et al., 1999; Gezer et al., 2015). In this case, too, the photometric period represents the orbital period. RVb systems, in fact, are often regarded as CE progenitors (see for example Kiss & Bódi 2017). Binaries with similar periods and eccentricity have also been detected in post-AGB systems (van Winckel et al., 2009; Oomen et al., 2018).
However, the above scenarios individually fail to account for the anti-correlated behavior in the and the redder bands. A combination of both effects, however, could explain the variability. In an eccentric orbit, interaction (either directly between the stars or with a circumstellar/binary disc) can be enhanced near periastron. This can lead to the significant production of dust, leading to increased extinction in the blue bands. In the red bands, both irradiation of the cooler component and enhanced dust emission can contribute to the brightening. The stochastic nature of the interaction may also explain any offset between the peaks in the red and the dip in the blue. Note that anticorrelation between ZTF- and bands were also recently reported in another system (PM 1-322, Paunzen et al. 2023, now classified to be a symbiotic candidate on HASH), and a similar scenario was proposed as a possibility.
We now briefly consider possible cases where the photometric period does not relate to binarity. Rotational modulation is often a cause for periodic variability, but, in most cases, the light curves in different bands are in phase. C25 compares it with light curves of a group of magnetic and chemically peculiar stars called Canum Venaticorum (ACV), which do indeed show anti-correlation between blue and red bands (see for example Gröbel et al. 2017; Faltová et al. 2021). However, the periods of these stars are much shorter (all known objects having days, Faltová et al. 2021) compared to that inferred in NGC 6833. On a different possibility, broadly similar light curves are witnessed in systems involving Be stars losing mass periodically due to pulsational instabilities. The lost mass appears as cold dust emitting mostly in far optical to mid-IR, leading to the brightening. Thus, it might be possible for the dust to induce a dip in the blue bands through extinction. In most such cases however, the ingress time is significantly shorter than the egress time and the infrared amplitudes are much larger than observed in NGC 6833 (see, for example, the light curves presented in Froebrich et al. 2023). This makes this scenario, too, unlikely.
All the above scenarios are expected to result mostly in the variability of the stellar continuum. However, any variation of the very strong emission lines can also induce significant photometric variability. We use the spectrum to check if the observed variability may result solely from emission line variations. We consider the ZTF- band () which shows the largest amplitude of . We calculate the emission line fluxes by appropriately modeling and subtracting the continuum. The emission line (including the Paschen series) contribution to the total flux is . With this spectrum taken at the -band maximum, this would require a significant decrement in the line strengths of all the emission lines to contribute meaningfully to the variability, which is unlikely. This, however, may not hold for the other bands, especially where the emissions are strong. In the ZTF- and bands, for example, the emission line comprises and of the total flux. Additionally, a significant fraction of the continuum is expected to be of nebular origin. Further epochal spectroscopic observations and proper modeling of the nebular continuum are needed to identify the source of the variability.
We briefly mention here that, in case the photometric variation is of stellar origin, the strong emission lines (and possibly the nebular continuum) can significantly suppress the true variability of the CSPN. We test this by assuming that the variability arises solely from the continuum. We calibrate the ZTF flux to the DBSP spectrum (in the respective bands) followed by subtraction of the emission line fluxes to get revised variability amplitudes. The amplitude of the dip in the increased to . The peaks in the and bands increased to relative flux values of and respectively. Thus, the band may not have an intrinsically higher amplitude than the band, as appears in the raw light curve. Though this exercise is preliminary and can only yield crude estimates, it shows that the effect can be significant. Interpretations, thus, need to be made with sufficient care.
3.2 NGC 6905 and Kn 26
NGC 6905 is a spectroscopically well-studied planetary nebula. The central star of NGC 6905, HD 193949, is a [WR]-type star with a temperature of kK (Gómez-González et al., 2022). The spectral classification seems to be uncertain, with reports of it being both [WO] (for example see Gómez-González et al. 2022) and [WC] (for example see Córsico et al. 2021) existing in the literature. The star also shows pulsations with periods in the range of seconds (Ciardullo & Bond, 1996; Córsico et al., 2021). The nebula shows a bipolar morphology. Kn 26, on the other hand, is a known quadrupolar PN (Guerrero et al., 2013), with Lanning 384 being the central star. Guerrero et al. (2013) finds the blue-side SED to be consistent with a blackbody of kK. Unlike the other, Kn 26 has not previously been studied spectroscopically.
3.2.1 Light Curves
Both NGC 6905 and Kn 26 show long-term modulation in the ZTF light curve. The Lomb-Scargle periodogram over the long-period range (same as mentioned in caption of Figure 2) shows a very prominent peak at days and a weaker peak at days for NGC 6905. C25 reports the longer period in their work. For Kn 26, two peaks of similar powers at days and days (the former being slightly dominant) appear in the periodogram444C25 does not report the long-period for Kn 26. One might guess that this is due to the use of standard photometry against difference photometry. However, the long periods in both the objects are also recovered from the standard photometry data. This in turn shows that the photometric behavior is not an artifact of difference imaging.. The periods appear individually in the and bands (see Figures 4 and 5, top panel). It is interesting to note that the inferred periods in the two objects are quite similar. This, and the closeness of the periods to possible systematic timescales (seasonal breaks and yearly observation cycles) prompted a few tests on the reality of the variability/periodicity. The details of the tests (which includes investigating the ATLAS data) can be found in Appendix B. We find no reason to reject the periods. Continued photometric monitoring is needed to confirm both the periodicity and the significance of the two peaks in the periodogram.
In addition to the long period, Kn 26 also shows a short-period signal at hours. When a Lomb-Scargle analysis over the short-period range (200 seconds to 100 days, same as used in Paper I) is applied on the raw light curve, the corresponding peak is visible, but it is subdominant to the -day cadence peak. However, the peak appears dominantly when the long-period modulation is subtracted out from the light curve and the periodogram is re-run. This result is presented in the third panel in Figure 5. The period is independently inferred in both the ZTF bands.
We used follow-up photometric observations to verify this period. Kn 26 was observed with the Wide Field Camera (WFC) at the Isaac Newton Telescope (INT) in July, 2021 in the -band with s exposure, and in August, 2022 in - and - bands with s and s exposures, respectively. The data were debiased and flat-fielded, before performing differential photometry of the central stars against non-variable field stars, all using standard astropy routines (Astropy Collaboration et al., 2018). Recently, we re-observed Kn 26 with the Caltech HIgh-speed Multi-color camERA (CHIMERA) in and bands on September, 2024. The CHIMERA images were corrected using bias and flat images555https://github.com/caltech-chimera/PyChimera. Photometry was extracted using aperture photometry (with a variable aperture size) using a modified version of the ULTRACAM pipeline (Feline et al., 2005). We made a differential light curve using a single reference star. GPS timestamps were used to determine the mid-exposure time of each image.
3.2.2 Follow-up Optical Spectroscopy
As noted earlier, the CSPN of NGC 6905 is spectroscopically very well studied. This is not the case with Kn 26, which prompted us to perform a spectroscopic observation. We took four 900 s exposures of Kn 26 at the twin 8.4m Large Binocular Telescope (LBT) using the Multi-Object Double Spectrographs (MODS, Pogge et al. 2010) on December, 2024. MODS provides two-channel grating spectroscopy by using a dichroic that splits the light at Å into separately optimized red and blue channels. The spectra cover the wavelength region Å and Å with a resolving power of and , respectively. The spectra were reduced using the modsccdred666https://github.com/rwpogge/modsCCDRed PYTHON package (Pogge, 2019) for basic 2d CCD reductions, and the modsidl777https://github.com/rwpogge/modsIDL pipeline (Croxall & Pogge, 2019) to extract 1d spectra and apply wavelength and flux calibrations.
The coadded MODS spectrum is shown in Fig. 6. We note that the Balmer and He II lines appear to be a super composition of a narrow emission line (likely resulting from the PN) and weaker but broader emission line (possibly originating from the photosphere). We identify the N III C III C IV emission line complex around 4650 Å but note that the the C IV 5801, 5812 Å doublet is not visible. Therefore, Kn 26 can be considered as a weak emission-line type ([WELS]) central star (Marcolino & de Araújo, 2003). We note here that the emission lines can also arise from the irradiated companion (as often the case with WELS type PN nuclei, see for example Miszalski et al. 2011a), and can be an indirect confirmation of the binarity of the system.
3.2.3 Discussion
The possible reasons behind the long-period variability observed in both the objects are, at this stage, uncertain. Binarity often results in long-term photometric modulation though ellipsoidal variability and/or irradiation effect. The spectrum of Kn 26, for example, might speak in favor of the latter as the weak emission lines may originate from an irradiated companion. Other causes of photometric variability is also possible, for example through stellar surface activity phased with the orbit, as suspected for the wide-binary LoTr 5 (see Kővári et al. 2019). In all of these cases, the photometric period is directly related to the orbital period (equal or double). Thus, both NGC 6905 and Kn 26 stand as candidates for long-period (maybe post-CE) wide binary systems. On a different possibility, long-timescale quasi-periodic modulation may result from close-binary involving a magnetically active giant companion: RS CVn systems. However, the associated long-periods are usually much longer ( years, see Martínez et al. 2022) than seen in NGC 6905 and Kn 26.
The light curves also bear a resemblance to another class of phenomena: Long Secondary Period (LSP) photometric modulations observed in several RGB and post-AGB stars (for example, see Kiss et al. 2006, Nicholls et al. 2009, and Soszyński et al. 2021). The phenomenon is not yet well understood and several possibilities have been proposed over the past couple of decades. One is related to binarity, where the companion is a low-mass object (large planet or brown dwarf), which carries a dust cloud around it, causing the occultation of the star (Soszyński et al., 2021). In such a scenario, we would expect an infrared excess in the spectral energy distribution (SED) of the CSPN.888We retrieved the SED of CTSS 2 and K 3-5 (around ′′ from the target coordinate) from the VizierR online photometry viewer tool http://vizier.cds.unistra.fr/vizier/sed/. We rejected all data points for which no errors are quoted. The SEDs for CTSS 2 and K 3-5 (discussed in Section 3.3) have been retrieved in the same way. Interestingly, strong infrared excess is seen in both NGC 6905 and Kn 26 (Figure 7). However, this excess is unlikely to be caused by a localized dust cloud around a companion, as the dimensions of the cloud required to match the observed excess are large. For example, the mid- and far-infrared excess in NHC 6905 is matched with a blackbody of temperature of K and radius of AU. This still does not rule out this possibility as the dust cloud causing the photometric variation might just be a smaller embedded component.
Another explanation for LSPs is pulsations of RGB and AGB stars (see for example Wood et al. 2004). It is not straightforward to extend this interpretation to CSPNe, as the white dwarfs are too compact to oscillate at such long periods. Their oscillation periods are usually much shorter (on the order of minutes, as in case of NGC 6905). This may again indicate binarity with a pulsating RGB and AGB companion. This is unlikely but definitely still a possibility.
@@ 3.2.3 Discussion
Pertaining to the short period, an explanation can be a close binary. The phase folded light curve is characterized by “pointed” minimas, morphologically similar to ellipsoidal variability with grazing eclipses. In that case, the orbital period is 2.36 hr, still making it the candidate for the shortest known binary period. In that case it would either be a double degenerate or a white dwarf – subdwarf system. Assuming typical masses of these stars yields a orbital semi-major axis of (consistent with ellipsoidal variability) and velocity (of each star) to be more than . This should be detectable if the inclination is sufficiently high. However, given that we do not see deep eclipses might indicate a low inclination, and thus, a low percieved radial velocity. The LBT spectra at different epochs do not show any obvious radial velocity shift. Thus this scenario too remains to be confirmed. But, if the short period indeed signifies a binarity, the interpretation of the long period becomes even more interesting and might indicate the possibility of a triple system inside a planetary nebula.
3.3 CTSS 2 and PN K 3-5
@@
CTSS 2 (PN G044.1+05.8) is a True PN on HASH. The nebula is faint and can be seen only in deep narrow-band images (see Figure 12.8 in Stanghellini et al. 2016 for the HST image). There has been some uncertainty in the literature regarding the true nature of the central star of CTSS 2. Stanghellini et al. (2016) notes that the identified CSPN appears too bright to be associated with the faint and evolved PN, and considers the possibility of it being an unrelated field star. They also consider the possibility of it being a binary, as does Moreno-Ibáñez et al. (2016). On the other hand, García-Hernández & Górny (2014) found emission signatures unusual for a PN in the Spitzer IR spectrum, and identified CTSS 2 to be a symbiotic system candidate. Confirmatory signatures of it being either a PN or symbiotic system are, however, lacking.
@@
PN K 3-5 (PN G034.3+06.2) is also a True PNe on HASH. It is a resolved bipolar nebula, with an angular diameter of . The most recent spectroscopic study was performed in the spectral survey of central stars by Weidmann et al. (2018). They indicated the presence of emission lines both from the central stellar region and the extended nebula, thus making it a candidate double-envelope PN.
@@ 3.3.2 Follow-up Optical Spectroscopy
CTSS 2 and K 3-5 were observed with the Low-Resolution Spectrograph (LRS2) on the 10-m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Hill et al. 2021) located at the McDonald Observatory in West Texas, USA, as a part of a survey for faint PN central stars (undertaken by H.E.B, see Bond et al. 2023 and subsequent papers). CTSS 2 was observed four times: twice on 15th September, 2024 with 250 seconds exposure each and twice again on 15th October, 2024 with 300 seconds exposure each. K 3-5 was observed once on 10th August, 2024 with 900 seconds exposure.
The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.999https://github.com/grzeimann/Panacea,101010https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.).
The narrow-band (NB) images for both objects around the brightest emission lines in the spectrum: H and [O III] 4363, 4959, 5007 Å are shown in Figure 9. From the two-dimensional images, one-dimensional spectra were extracted for both the central core (nucleus) and the nebula. The corresponding apertures are shown within the NB images (white and black dashed apertures correspond to the nucleus and nebula, respectively). The extracted spectra for CTSS 2 and K 3-5 are shown in Figures 10 (top panel) and 11 respectively.
Both CTSS 2 and K 3-5 were also observed with DBSP on 13th May, 2024 with exposures of 15 minutes. It is different from the HET spectrum only for its wider coverage in the red arm. However, the HET spectrum is of higher quality and, being an IFU, enables more accurate analyses. Thus, we base our discussion on the HET spectrum in the following sections. We provide the DBSP spectrum separately in Appendix C.

### v3

@@ 3.2.1 Light Curves
Both NGC 6905 and Kn 26 show long-term modulation in the ZTF light curve. The Lomb-Scargle periodogram over the long-period range (same as mentioned in caption of Figure 2) shows a very prominent peak at days and a weaker peak at days for NGC 6905. For Kn 26, two peaks of similar powers at days and days (the former being slightly dominant) appear in the periodogram. The flux semi-amplitude corresponding to the most prominent periods for the two objects are and , respectively. 666C25 reports the 700 day period for NGC 6905 in their work. They, however, do not report these long periods in Kn 26. One might guess that this is due to the use of standard photometry against difference photometry. However, the long periods in both the objects are also recovered from the standard photometry data. This in turn shows that the photometric behavior is not an artifact of difference imaging. The periods appear individually in the and bands (see Figures 3 and 4, top panel). It is interesting to note that the inferred periods in the two objects are quite similar. This, and the closeness of the periods to possible systematic timescales (seasonal breaks and yearly observation cycles) prompted several tests on the reality of the variability/periodicity. The details of the tests can be found in Appendix D. We find no reason to reject the periods.
In addition to the long periods, Kn 26 shows a short-period signal at hours. When a Lomb-Scargle analysis over the short-period range (200 seconds to 100 days, same as used in Paper I) is applied on the raw ZTF light curve, the corresponding peak is visible, but it is subdominant to the -day cadence peak. However, the peak appears dominantly when the long-period modulation is subtracted out from the light curve and the periodogram is re-run. This result is presented in the third panel in Figure 4. The period is independently inferred in both the ZTF bands.
We used follow-up photometric observations to verify this period. Kn 26 was observed with the Wide Field Camera (WFC) at the Isaac Newton Telescope (INT) in 2021 July in the -band with s exposure, and in 2022 August in - and -bands with s and s exposures, respectively. The data were debiased and flat-fielded, before performing differential photometry of the central stars against non-variable field stars, all using standard astropy routines (Astropy Collaboration et al., 2018). Recently, we re-observed Kn 26 with the Caltech HIgh-speed Multi-color camERA (CHIMERA) in and bands in 2024 September. The CHIMERA images were corrected using bias and flat images.777https://github.com/caltech-chimera/PyChimera Photometry was extracted using aperture photometry (with a variable aperture size) using a modified version of the ULTRACAM pipeline (Feline et al., 2005). We made a differential light curve using a single reference star. GPS timestamps were used to determine the mid-exposure time of each image.
@@ 3.2.2 Discussion
Ellipsoidal modulation or irradiation effect in a binary can result in the long-period variability. The systems can also be similar to LoTr 5, where the long-term photometric modulation is suspected to arise from stellar surface activity of a giant companion phased with the orbit (Kővári et al., 2019). This is reinforced by the fact that the variability amplitudes of our objects are similar (slightly larger) to that of LoTr 5. In all of these cases, the orbital period is directly related to the photometric period (equal or double). These make both NGC 6905 and Kn 26 additional candidates for long-period wide binary CSPNe.
The light curves also bear a resemblance to Long Secondary Period (LSP) photometric modulations observed in several RGB and post-AGB stars (for example, see Kiss et al. 2006, Nicholls et al. 2009, and Soszyński et al. 2021, though the amplitudes are usually larger than seen in our targets.). This phenomenon is not yet well understood. But wide binarity and phase-dependent dust occultation (similar to RVb systems) is one proposed explanation. Here, too, the period corresponds to the binary orbital period.
We now discuss two possibilities where the long period does not correspond to the orbital period, still requiring binarity. Magnetic activity cycles of giant stars in close binaries (RS CVn systems) often result in long quasi-periodic photometric modulation. However, the associated periods are usually much longer ( years, see Martínez et al. 2022) than seen in NGC 6905 and Kn 26. Thus, this scenario is unlikely. The other possibility is stellar pulsations of giant stars, which is another widely accepted explanations for LSPs. Pawlak (2021) showed that LSPs tend to be stars in the RGB or AGB phase transitioning between period-luminosity sequences. Thus, with a twin-binary (El-Badry et al., 2019) setup it is possible that the companion of the CSPN is a pulsating LSP. High resolution radial velocity measurements is required to distinguish between the different scenarios.
@@
The short period of 1.18 hours can arise from rotational modulation of the white dwarf (Oliveira da Rosa et al., 2024). Another more exotic, but unlikely, possibility (also mentioned in C25) is a close binary (this, with the long periods, may indicate a triple system!). The phase folded light curve is characterized by “pointed” minimas, morphologically similar to ellipsoidal variability with grazing eclipses. In that case, the orbital period is 2.36 hr, making it the candidate for the shortest known CSPN binary period. At such short periods, for typical stellar parameters, we expect a large radial velocity variation of . This should be detectable. The LBT spectra at different epochs, however, do not show any radial velocity shift.
3.3 CTSS 2 and K 3-5
@@
CTSS 2 (PN G044.1+05.8) is a True PN in the HASH catalog. The nebula is faint and can be seen only in deep narrow-band images (see Figure 12.8 in Stanghellini et al. 2016 for the HST image). There has been some uncertainty in the literature regarding the true nature of the central star of CTSS 2. Stanghellini et al. (2016) note that the identified CSPN appears too bright to be associated with the faint and evolved PN, and considers the possibility of it being an unrelated field star. They also consider the possibility of it being a binary, as do Moreno-Ibáñez et al. (2016). On the other hand, García-Hernández & Górny (2014) found emission signatures unusual for a PN in the Spitzer IR spectrum, and identified CTSS 2 to be a symbiotic system candidate. Confirmatory signatures of it being either a PN or symbiotic system are, however, lacking.
@@
K 3-5 (PN G034.3+06.2) is also a True PNe in HASH. It is a resolved bipolar nebula, with an angular diameter of . The most recent spectroscopic study was performed in the spectral survey of central stars by Weidmann et al. (2018). They indicated the presence of emission lines both from the central stellar region and the extended nebula, thus making it a candidate double-envelope PN.
@@ 3.3.2 Follow-up Optical Spectroscopy
CTSS 2 and K 3-5 were observed with the Low-Resolution Spectrograph (LRS2) on the 10-m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Hill et al. 2021) located at the McDonald Observatory in west Texas, USA, as a part of a spectroscopic survey of faint PN central stars (undertaken by H.E.B. and collaborators; see Bond et al. 2023 and subsequent papers). CTSS 2 was observed twice: 2024 September 15, s, and 2024 October 15, s. K 3-5 was observed once, on 2024 August 10, with a 900 s exposure.
The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.888https://github.com/grzeimann/Panacea,999https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.).
The narrow-band (NB) images for both objects around the brightest emission lines in the spectrum: H and [O III] 4363, 4959, 5007 Å are shown in Figure 6. From the two-dimensional images, one-dimensional spectra were extracted for both the central core (nucleus) and the nebula. The corresponding apertures are shown within the NB images (white and black dashed apertures correspond to the nucleus and nebula, respectively). The extracted spectra for CTSS 2 and K 3-5 are shown in Figures 7 (top panel) and 8 respectively.
Both CTSS 2 and K 3-5 were also observed with DBSP on 13th May, 2024 with exposures of 15 minutes. It is different from the HET spectrum only for its wider coverage in the red arm. However, the HET spectrum is of higher quality and, being an IFU, enables more accurate analyses. Thus, we base our discussion on the HET spectrum in the following sections. We provide the DBSP spectrum separately in Appendix E.